OpenStudy (anonymous):

Find the domain of the exponential function f(x) = (1/3)x. A. Any real numbers B. Positive real numbers C. Negative real numbers D. Non-zero real numbers Find the range of the exponential function f(x) = (1/3)x. A. Any real numbers B. Positive real numbers C. Negative real numbers D. Non-zero real numbers last ones

OpenStudy (anonymous):

1.) f(x) = (1/3)^x 2.) f(x) = (1/3)^x